# Latent Factor Analysis in Short Panels

### v2

@@ E.1 Pseudo likelihood and PML estimator
E.2 Global identification and consistency
The population criterion $L_{0}(\theta)$ is defined in Appendix A, with $V_{y}=V_{y}^{0}=\Sigma(\theta_{0})=F_{0}F_{0}^{\prime}+V_{\varepsilon}^{0}$.
@@ E.2 Global identification and consistency
The following conditions are equivalent: a) the true value $\theta_{0}$ is the unique maximizer of $L_{0}(\theta)$ for $\theta\in\Theta$; b) $\Sigma(\theta)=\Sigma(\theta_{0})$, $\theta\in\Theta$Rightarrow $\theta=\theta_{0}$, up to sign changes in the columns of $F$. They yield the global identification in the FA model.
In Lemma 4, condition a) is the standard identification condition for a M-estimator with population criterion $L_{0}(\theta)$. Condition (b) is the global identification condition based on the variance matrix as in Anderson and Rubin (1956). Condition (b) corresponds to our Assumption A.4.
Let us now establish the consistency of the FA estimators in our setting. Write $\hat{V}_{y}=\frac{1}{n}\sum_{i=1}^{n}(\varepsilon_{i}-\bar{\varepsilon})(\varepsilon_{i}-\bar{\varepsilon})^{\prime}+F[\frac{1}{n}\sum_{i=1}^{n}(\beta_{i}-\bar{\beta})(\beta_{i}-\bar{\beta})^{\prime}]F^{\prime}+F[\frac{1}{n}\sum_{i=1}^{n}(\beta_{i}-\bar{\beta})(\varepsilon_{i}-\bar{\varepsilon})^{\prime}]+[\frac{1}{n}\sum_{i=1}^{n}(\varepsilon_{i}-\bar{\varepsilon})(\beta_{i}-\bar{\beta})^{\prime}]F^{\prime}$, where $\bar{\varepsilon}=\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}$ and $\bar{\beta}=\frac{1}{n}\sum_{i=1}^{n}\beta_{i}$. Under the normalization in Assumption A.1 we have:
$\hat{V}_{y}=\frac{1}{n}\varepsilon\varepsilon^{\prime}-\bar{\varepsilon}\bar{\varepsilon}^{\prime}+FF^{\prime}+F\left(\frac{1}{n}\varepsilon\beta\right)^{\prime}+\left(\frac{1}{n}\varepsilon\beta\right)F^{\prime}.$ (E.2)
@@
[…]
@@
Under Assumptions 1, 2, and A.2-A.4, the FA estimators $\hat{F}$, $\hat{V}_{\varepsilon}$ and $\hat{\mu}$ are consistent as $n\rightarrow\infty$ and $T$ is fixed.
@@
E.3 Asymptotic expansions of estimators $\hat{V}_{\varepsilon}$ and $\hat{F}$
The FA estimators $\hat{V}_{\varepsilon}$ and $\hat{F}$ are consistent M-estimators under nonlinear constraints, and admit expansions at first order for fixed $T$ and $n\rightarrow\infty$, namely $\hat{V}_{\varepsilon}=\tilde{V}_{\varepsilon}+\frac{1}{\sqrt{n}}\Psi_{\varepsilon}+o_{p}(\frac{1}{\sqrt{n}})$ and $\hat{F}_{j}=F_{j}+\frac{1}{\sqrt{n}}\Psi_{F_{j}}+o_{p}(\frac{1}{\sqrt{n}})$ (see Appendix E.5.1). The next proposition (new to the literature) characterizes the diagonal random matrix $\Psi_{\varepsilon}$ and the random vectors $\Psi_{F_{j}}$ by using conditions (FA1) and (FA2) in Section 2 (see proof at the end of the section).
@@ E.3 Asymptotic expansions of estimators $\hat{V}_{\varepsilon}$ and $\hat{F}$
Under Assumptions 1, 2, and A.1-A.4, A.6, we have (a) for $j=1,...,k$
$\Psi_{F_{j}}$ $=$ $R_{j}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F_{j}+\Lambda_{j}\Psi_{\varepsilon}V_{\varepsilon}^{-1}F_{j},$ (E.3)
@@
$diag\left(M_{F,{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})M_{F,{V}_{\varepsilon}}^{\prime}\right)=0.$ (E.4)
@@
Equation (E.3) yields the asymptotic expansion of the eigenvectors by accounting for estimation errors of matrix $\hat{V}_{y}\hat{V}_{\varepsilon}^{-1}$ (first term) and of the normalization constraint (second term). To interpret Equation (E.4), we can observe that the matrix $M_{F,{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})M_{F,{V}_{\varepsilon}}^{\prime}$ yields the first-order term in the asymptotic expansion of $\sqrt{n}\hat{S}$ (up to the left- and right-multiplication by diagonal matrix $V_{\varepsilon}^{-1/2}$). Thus, Equation (E.4) is implied by the property that the diagonal terms of matrix $\hat{S}$ are equal to zero as stated in Proposition 1 (c).
@@
Let us now give the explicit expression of $\Psi_{\varepsilon}$. By using $M_{F,{V}_{\varepsilon}}\Psi_{y}M_{F,{V}_{\varepsilon}}^{\prime}=M_{F,{V}_{\varepsilon}}Z_{n}M_{F,{V}_{\varepsilon}}^{\prime}$, we can rewrite Equation (E.4) as $diag\left(M_{F,{V}_{\varepsilon}}(Z_{n}-\Psi_{\varepsilon})M_{F,{V}_{\varepsilon}}^{\prime}\right)=0$. Now, since $\Psi_{\varepsilon}$ is diagonal, we have $diag\left(M_{F,{V}_{\varepsilon}}\Psi_{\varepsilon}M_{F,{V}_{\varepsilon}}^{\prime}\right)=M_{F,{V}_{\varepsilon}}^{\odot 2}diag(\Psi_{\varepsilon})$, where $M_{F,{V}_{\varepsilon}}^{\odot 2}=M_{F,{V}_{\varepsilon}}\odot M_{F,{V}_{\varepsilon}}$. Thus, we get:
$M_{F,{V}_{\varepsilon}}^{\odot 2}diag(\Psi_{\varepsilon})=diag(M_{F,{V}_{\varepsilon}}Z_{n}M_{F,{V}_{\varepsilon}}^{\prime}).$ (E.5)
@@
To have a unique solution for vector $diag(\Psi_{\varepsilon})$, we need the non-singularity of the $T\times T$ matrix $M_{F,{V}_{\varepsilon}}^{\odot 2}$. It is the local identification condition in the FA model stated in Assumption A.5. Let us write $G=[g_{1}:\cdots:g_{T-k}]$. Then, we have $M_{F,V_{\varepsilon}}=GG^{\prime}V_{\varepsilon}^{-1}=\sum_{j=1}^{T-k}g_{j}(V_{\varepsilon}^{-1}g_{j})^{\prime}$, and so we get the Hadamard product $M_{F,V_{\varepsilon}}^{\odot 2}=\sum_{i,j=1}^{T-k}[g_{i}(V_{\varepsilon}^{-1}g_{i})^{\prime}]\odot[g_{j}(V_{\varepsilon}^{-1}g_{j})^{\prime}]=\left[\sum_{i,j=1}^{T-k}(g_{i}\odot g_{j})(g_{i}\odot g_{j})^{\prime}\right]$ $V_{\varepsilon}^{-2}$ $=2\left(\boldsymbol{X}^{\prime}\boldsymbol{X}\right)V_{\varepsilon}^{-2}.$ [^36] Let us recall the following property of the Hadamard product: $(ab^{\prime})\odot(cd^{\prime})=(a\odot c)(b\odot d)^{\prime}$ for conformable vectors $a,b,c,d$. The last equalitiy because $\boldsymbol{X}^{\prime}=\left[\frac{1}{\sqrt{2}}g_{1}\odot g_{1}~{}:~{}\cdots~{}:~{}\frac{1}{\sqrt{2}}g_{T-k}\odot g_{T-k}~{}:~{}\{g_{i}\odot g_{j}\}_{i<j}\right]$ (see beginning of the proof of Proposition 2 (a)) . Hence, we can state the local identification condition in Assumption A.5 as a full-rank condition for matrix $\boldsymbol{X}$, analogously as in linear regression (Lemma 6). In Lemma 6 in Appendix E.4 i), we also show equivalence with invertibility of the bordered Hessian, i.e., the Hessian of the Lagrangian function in a constrained M-estimation.
Under Assumption A.5, we get from Equation (E.5):
$\Psi_{\varepsilon}=\mathcal{T}_{F,V_{\varepsilon}}(Z_{n}),$ (E.6)
@@
$diag(\Psi_{\varepsilon})=V_{\varepsilon}^{2}\left(\boldsymbol{X}^{\prime}\boldsymbol{X}\right)^{-1}\boldsymbol{X}^{\prime}vech\left(Z^{*}_{n}\right).$ (E.7)
@@
Anderson and Rubin (1956), Theorem 12.1, show that the FA estimator is asymptotically normal if $\sqrt{n}(\hat{V}_{y}-V_{y})$ is asymptotically normal. They use a linearization of the first-order conditions similar as the one of Proposition 7. Their Equation (12.16) corresponds to our Equation (E.4). However, they only provide an implicit characterization of the $\Psi_{F_{j}}$ and not an explicit expression for $\Psi_{\varepsilon}$ and $\Psi_{F_{j}}$ in terms of asymptotically Gaussian random matrices like $Z_{n}$ as we do. These key developments pave the way to establishing the asymptotic distributions of estimators $\hat{F}$ and $\hat{V}_{\varepsilon}$ in general settings, that we cover in Appendix E.5.
@@
Proof of Proposition 7: From (E.2) and Lemma 5 we have $\hat{V}_{y}=\tilde{V}_{y}+\frac{1}{\sqrt{n}}\Psi_{y}+o_{p}(\frac{1}{\sqrt{n}})$, where $\tilde{V}_{y}=FF^{\prime}+\tilde{V}_{\varepsilon}$ and $\Psi_{y}=\frac{1}{\sqrt{n}}(\varepsilon\beta F^{\prime}+F\beta^{\prime}\varepsilon^{\prime})+\sqrt{n}\left(\frac{1}{n}\varepsilon\varepsilon^{\prime}-\tilde{V}_{\varepsilon}\right)$. Let us substitute this expansion for $\hat{V}_{y}$ into (FA2) and rearrange to obtain $\hat{F}\hat{\Gamma}-FF^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F}=\frac{1}{\sqrt{n}}\Psi_{y}\hat{V}_{\varepsilon}^{-1}\hat{F}+(\tilde{V}_{\varepsilon}\hat{V}_{\varepsilon}^{-1}-I_{T})\hat{F}+o_{p}(\frac{1}{\sqrt{n}})$, where $\hat{\Gamma}=\hat{F}^{\prime}\hat{V}^{-1}_{\varepsilon}\hat{F}=diag(\hat{\gamma}_{1},\dots,\hat{\gamma}_{k})$. From $\hat{V}_{\varepsilon}=\tilde{V}_{\varepsilon}+\frac{1}{\sqrt{n}}\Psi_{\varepsilon}+o_{p}(\frac{1}{\sqrt{n}})$, we have $\tilde{V}_{\varepsilon}\hat{V}_{\varepsilon}^{-1}-I_{T}=-\frac{1}{\sqrt{n}}\Psi_{\varepsilon}\hat{V}_{\varepsilon}^{-1}+o_{p}(\frac{1}{\sqrt{n}})$. Substituting into the above equation and right multiplying both sides by $(F^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F})^{-1}$ gives $\hat{F}\hat{\mathcal{D}}-F=\frac{1}{\sqrt{n}}(\Psi_{y}-\Psi_{\varepsilon})\hat{V}_{\varepsilon}^{-1}\hat{F}(F^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F})^{-1}+o_{p}(\frac{1}{\sqrt{n}})$, where $\hat{\mathcal{D}}:=\hat{\Gamma}(F^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F})^{-1}$. By the root-$n$ convergence of the FA estimates (see Section E.5.1), we get
$\hat{F}\hat{\mathcal{D}}-F=\frac{1}{\sqrt{n}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+o_{p}(\frac{1}{\sqrt{n}}),$ (E.8)
@@
and $\hat{\mathcal{D}}=I_{k}+O_{p}(\frac{1}{\sqrt{n}})$, where $\Gamma=diag(\gamma_{1},...,\gamma_{k})$. We can push the expansion by plugging into (E.8) the expansion of $\hat{\mathcal{D}}$. We have $F^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F}=[I_{k}-(\hat{F}-F)^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F}\hat{\Gamma}^{-1}]\hat{\Gamma}$, so that $\hat{\mathcal{D}}=[I_{k}-(\hat{F}-F)^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F}\hat{\Gamma}^{-1}]^{-1}=I_{k}+(\hat{F}-F)^{\prime}V_{\varepsilon}^{-1}F\Gamma^{-1}+o_{p}(\frac{1}{\sqrt{n}})$. By plugging into (E.8), we get:
$\hat{F}-F+F[(\hat{F}-F)^{\prime}V_{\varepsilon}^{-1}F\Gamma^{-1}]=\frac{1}{\sqrt{n}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+o_{p}(\frac{1}{\sqrt{n}}).$ (E.9)
@@
By multiplying both sides with $M_{F,V_{\varepsilon}}$, we get $M_{F,V_{\varepsilon}}(\hat{F}-F)=\frac{1}{\sqrt{n}}M_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+o_{p}(\frac{1}{\sqrt{n}})$. Then, $\hat{F}-F=\frac{1}{\sqrt{n}}M_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+\frac{1}{\sqrt{n}}FA+o_{p}(\frac{1}{\sqrt{n}})$, where $A$ is a random $k\times k$ matrix to be determined next. By plugging into (E.9), we get $F(A+A^{\prime})=P_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+o_{p}(\frac{1}{\sqrt{n}})$. By multiplying both sides by $\frac{1}{2}\Gamma^{-1}F^{\prime}V_{\varepsilon}^{-1}$ and using $F^{\prime}V_{\varepsilon}^{-1}P_{F,V_{\varepsilon}}=F^{\prime}V_{\varepsilon}^{-1}$, we get the symmetric part of matrix $A$, i.e., $\frac{1}{2}(A+A^{\prime})=\frac{1}{2}\Gamma^{-1}F^{\prime}V_{\varepsilon}^{-1}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}$ (we include higher-order terms in the remainder $o_{p}(\frac{1}{\sqrt{n}})$). Thus, $\hat{F}-F=\frac{1}{\sqrt{n}}\Psi_{F}+o_{p}(\frac{1}{\sqrt{n}})$, where
$\Psi_{F}=M_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+\frac{1}{2}P_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+F\tilde{A},$ (E.10)

### v3

@@ B.1 Pseudo likelihood and PML estimator
B.2 Global identification and consistency
The population criterion $L_{0}(\theta)$ is defined in Section 3, with $V_{y}=V_{y}^{0}=\Sigma(\theta_{0})=F_{0}F_{0}^{\prime}+V_{\varepsilon}^{0}$.
@@ B.2 Global identification and consistency
The following conditions are equivalent: a) the true value $\theta_{0}$ is the unique maximizer of $L_{0}(\theta)$ for $\theta\in\Theta$; b) $\Sigma(\theta)=\Sigma(\theta_{0})$, $\theta\in\Theta$ $\Rightarrow$ $\theta=\theta_{0}$, up to sign changes in the columns of $F$. They yield the global identification in the FA model.
In Lemma 5, condition a) is the standard identification condition for a M-estimator with population criterion $L_{0}(\theta)$. Condition (b) is the global identification condition based on the variance matrix as in Anderson and Rubin (1956). Condition (b) corresponds to our Assumption 6.
Let us now establish the consistency of the FA estimators in our setting. Write $\hat{V}_{y}=\frac{1}{n}\sum_{i=1}^{n}(\varepsilon_{i}-\bar{\varepsilon})(\varepsilon_{i}-\bar{\varepsilon})^{\prime}+F[\frac{1}{n}\sum_{i=1}^{n}(\beta_{i}-\bar{\beta})(\beta_{i}-\bar{\beta})^{\prime}]F^{\prime}+F[\frac{1}{n}\sum_{i=1}^{n}(\beta_{i}-\bar{\beta})(\varepsilon_{i}-\bar{\varepsilon})^{\prime}]+[\frac{1}{n}\sum_{i=1}^{n}(\varepsilon_{i}-\bar{\varepsilon})(\beta_{i}-\bar{\beta})^{\prime}]F^{\prime}$, where $\bar{\varepsilon}=\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}$ and $\bar{\beta}=\frac{1}{n}\sum_{i=1}^{n}\beta_{i}$. Under the normalization in Assumption 3 we have:
$\hat{V}_{y}=\frac{1}{n}\varepsilon\varepsilon^{\prime}-\bar{\varepsilon}\bar{\varepsilon}^{\prime}+FF^{\prime}+F\left(\frac{1}{n}\varepsilon\beta\right)^{\prime}+\left(\frac{1}{n}\varepsilon\beta\right)F^{\prime}.$ (B.2)
@@
Under Assumptions 1, 2, and 4, 5, as $n\rightarrow\infty$, we have: (a) $\bar{\varepsilon}=o_{p}(\frac{1}{n^{1/4}})$, (b) $\frac{1}{n}\varepsilon\varepsilon^{\prime}\overset{p}{\rightarrow}V_{\varepsilon}^{0}$, and (c) $\frac{1}{n}\varepsilon\beta\overset{p}{\rightarrow}0$.
From Equation (B.2) and Lemma 6, we have $\hat{V}_{y}\overset{p}{\rightarrow}V_{y}^{0}$. Thus, $\hat{L}(\theta)$ converges in probability to $L_{0}(\theta)$ as $n\rightarrow\infty$, uniformly over $\Theta$ compact. From standard results on M-estimators, we get consistency of $\hat{\theta}$. Moreover, from $\bar{y}=\mu+\bar{\varepsilon}$, we get the consistency of $\hat{\mu}$.
@@
Under Assumptions 1, 2, and 4-6, the FA estimators $\hat{F}$, $\hat{V}_{\varepsilon}$ and $\hat{\mu}$ are consistent as $n\rightarrow\infty$ and $T$ is fixed.
@@
B.3 Asymptotic expansions of estimators $\hat{V}_{\varepsilon}$ and $\hat{F}$
The FA estimators $\hat{V}_{\varepsilon}$ and $\hat{F}$ are consistent M-estimators under nonlinear constraints, and admit expansions at first order for fixed $T$ and $n\rightarrow\infty$, namely $\hat{V}_{\varepsilon}=\tilde{V}_{\varepsilon}+\frac{1}{\sqrt{n}}\Psi_{\varepsilon}+o_{p}(\frac{1}{\sqrt{n}})$ and $\hat{F}_{j}=F_{j}+\frac{1}{\sqrt{n}}\Psi_{F_{j}}+o_{p}(\frac{1}{\sqrt{n}})$ (see Appendix B.5.1). The next proposition (new to the literature) characterizes the diagonal random matrix $\Psi_{\varepsilon}$ and the random vectors $\Psi_{F_{j}}$ by using conditions (FA1) and (FA2) in Section 2 (see proof at the end of the section).
@@ B.3 Asymptotic expansions of estimators $\hat{V}_{\varepsilon}$ and $\hat{F}$
Under Assumptions 1, 2, and 3-6, 8, we have (a) for j=1,,kj=1,...,k
$\Psi_{F_{j}}$ $=$ $R_{j}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F_{j}+\Lambda_{j}\Psi_{\varepsilon}V_{\varepsilon}^{-1}F_{j},$ (B.3)
@@
$diag\left(M_{F,{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})M_{F,{V}_{\varepsilon}}^{\prime}\right)=0.$ (B.4)
@@
Equation (B.3) yields the asymptotic expansion of the eigenvectors by accounting for estimation errors of matrix $\hat{V}_{y}\hat{V}_{\varepsilon}^{-1}$ (first term) and of the normalization constraint (second term). To interpret Equation (B.4), we can observe that the matrix $M_{F,{V}_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})M_{F,{V}_{\varepsilon}}^{\prime}$ yields the first-order term in the asymptotic expansion of $\sqrt{n}\hat{S}$ (up to the left- and right-multiplication by diagonal matrix $V_{\varepsilon}^{-1/2}$). Thus, Equation (B.4) is implied by the property that the diagonal terms of matrix $\hat{S}$ are equal to zero as stated in Lemma 1 (c).
@@
Let us now give the explicit expression of $\Psi_{\varepsilon}$. By using $M_{F,{V}_{\varepsilon}}\Psi_{y}M_{F,{V}_{\varepsilon}}^{\prime}=M_{F,{V}_{\varepsilon}}Z_{n}M_{F,{V}_{\varepsilon}}^{\prime}$, we can rewrite Equation (B.4) as $diag\left(M_{F,{V}_{\varepsilon}}(Z_{n}-\Psi_{\varepsilon})M_{F,{V}_{\varepsilon}}^{\prime}\right)=0$. Now, since $\Psi_{\varepsilon}$ is diagonal, we have $diag\left(M_{F,{V}_{\varepsilon}}\Psi_{\varepsilon}M_{F,{V}_{\varepsilon}}^{\prime}\right)=M_{F,{V}_{\varepsilon}}^{\odot 2}diag(\Psi_{\varepsilon})$, where $M_{F,{V}_{\varepsilon}}^{\odot 2}=M_{F,{V}_{\varepsilon}}\odot M_{F,{V}_{\varepsilon}}$. Thus, we get:
$M_{F,{V}_{\varepsilon}}^{\odot 2}diag(\Psi_{\varepsilon})=diag(M_{F,{V}_{\varepsilon}}Z_{n}M_{F,{V}_{\varepsilon}}^{\prime}).$ (B.5)
@@
To have a unique solution for vector $diag(\Psi_{\varepsilon})$, we need the non-singularity of the $T\times T$ matrix $M_{F,{V}_{\varepsilon}}^{\odot 2}$. It is the local identification condition in the FA model stated in Assumption 7. Let us write $G=[g_{1}:\cdots:g_{T-k}]$. Then, we have $M_{F,V_{\varepsilon}}=GG^{\prime}V_{\varepsilon}^{-1}=\sum_{j=1}^{T-k}g_{j}(V_{\varepsilon}^{-1}g_{j})^{\prime}$, and so we get the Hadamard product $M_{F,V_{\varepsilon}}^{\odot 2}=\sum_{i,j=1}^{T-k}[g_{i}(V_{\varepsilon}^{-1}g_{i})^{\prime}]\odot[g_{j}(V_{\varepsilon}^{-1}g_{j})^{\prime}]=\left[\sum_{i,j=1}^{T-k}(g_{i}\odot g_{j})(g_{i}\odot g_{j})^{\prime}\right]$ $V_{\varepsilon}^{-2}$ $2\left(\boldsymbol{X}^{\prime}\boldsymbol{X}\right)V_{\varepsilon}^{-2}.$ [^9] Let us recall the following property of the Hadamard product: $(ab^{\prime})\odot(cd^{\prime})=(a\odot c)(b\odot d)^{\prime}$ for conformable vectors $a,b,c,d$. The last equalitiy because $\boldsymbol{X}^{\prime}=\left[\frac{1}{\sqrt{2}}g_{1}\odot g_{1}~:~\cdots~:~\frac{1}{\sqrt{2}}g_{T-k}\odot g_{T-k}~:~\{g_{i}\odot g_{j}\}_{i<j}\right]$ (see beginning of the proof of Proposition 2 (a)) . Hence, we can state the local identification condition in Assumption 7 as a full-rank condition for matrix $\boldsymbol{X}$, analogously as in linear regression (Lemma 7). In Lemma 7 in Appendix B.4 i), we also show equivalence with invertibility of the bordered Hessian, i.e., the Hessian of the Lagrangian function in a constrained M-estimation.
Under Assumption 7, we get from Equation (B.5):
$\Psi_{\varepsilon}=\mathcal{T}_{F,V_{\varepsilon}}(Z_{n}),$ (B.6)
@@
$diag(\Psi_{\varepsilon})=V_{\varepsilon}^{2}\left(\boldsymbol{X}^{\prime}\boldsymbol{X}\right)^{-1}\boldsymbol{X}^{\prime}vech\left(Z^{*}_{n}\right).$ (B.7)
@@
Anderson and Rubin (1956), Theorem 12.1, show that the FA estimator is asymptotically normal if $\sqrt{n}(\hat{V}_{y}-V_{y})$ is asymptotically normal. They use a linearization of the first-order conditions similar as the one of Proposition 6. Their Equation (12.16) corresponds to our Equation (B.4). However, they only provide an implicit characterization of the $\Psi_{F_{j}}$ and not an explicit expression for $\Psi_{\varepsilon}$ and $\Psi_{F_{j}}$ in terms of asymptotically Gaussian random matrices like $Z_{n}$ as we do. These key developments pave the way to establishing the asymptotic distributions of estimators $\hat{F}$ and $\hat{V}_{\varepsilon}$ in general settings, that we cover in Appendix B.5.
@@
Proof of Proposition 6: From (B.2) and Lemma 6 we have $\hat{V}_{y}=\tilde{V}_{y}+\frac{1}{\sqrt{n}}\Psi_{y}+o_{p}(\frac{1}{\sqrt{n}})$, where $\tilde{V}_{y}=FF^{\prime}+\tilde{V}_{\varepsilon}$ and $\Psi_{y}=\frac{1}{\sqrt{n}}(\varepsilon\beta F^{\prime}+F\beta^{\prime}\varepsilon^{\prime})+\sqrt{n}\left(\frac{1}{n}\varepsilon\varepsilon^{\prime}-\tilde{V}_{\varepsilon}\right)$. Let us substitute this expansion for $\hat{V}_{y}$ into (FA2) and rearrange to obtain $\hat{F}\hat{\Gamma}-FF^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F}=\frac{1}{\sqrt{n}}\Psi_{y}\hat{V}_{\varepsilon}^{-1}\hat{F}+(\tilde{V}_{\varepsilon}\hat{V}_{\varepsilon}^{-1}-I_{T})\hat{F}+o_{p}(\frac{1}{\sqrt{n}})$, where $\hat{\Gamma}=\hat{F}^{\prime}\hat{V}^{-1}_{\varepsilon}\hat{F}=diag(\hat{\gamma}_{1},\dots,\hat{\gamma}_{k})$. From $\hat{V}_{\varepsilon}=\tilde{V}_{\varepsilon}+\frac{1}{\sqrt{n}}\Psi_{\varepsilon}+o_{p}(\frac{1}{\sqrt{n}})$, we have $\tilde{V}_{\varepsilon}\hat{V}_{\varepsilon}^{-1}-I_{T}=-\frac{1}{\sqrt{n}}\Psi_{\varepsilon}\hat{V}_{\varepsilon}^{-1}+o_{p}(\frac{1}{\sqrt{n}})$. Substituting into the above equation and right multiplying both sides by $(F^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F})^{-1}$ gives $\hat{F}\hat{\mathcal{D}}-F=\frac{1}{\sqrt{n}}(\Psi_{y}-\Psi_{\varepsilon})\hat{V}_{\varepsilon}^{-1}\hat{F}(F^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F})^{-1}+o_{p}(\frac{1}{\sqrt{n}})$, where $\hat{\mathcal{D}}:=\hat{\Gamma}(F^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F})^{-1}$. By the root-$n$ convergence of the FA estimates (see Section B.5.1), we get
$\hat{F}\hat{\mathcal{D}}-F=\frac{1}{\sqrt{n}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+o_{p}(\frac{1}{\sqrt{n}}),$ (B.8)
@@
and $\hat{\mathcal{D}}=I_{k}+O_{p}(\frac{1}{\sqrt{n}})$, where $\Gamma=diag(\gamma_{1},...,\gamma_{k})$. We can push the expansion by plugging into (B.8) the expansion of $\hat{\mathcal{D}}$. We have $F^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F}=[I_{k}-(\hat{F}-F)^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F}\hat{\Gamma}^{-1}]\hat{\Gamma}$, so that $\hat{\mathcal{D}}=[I_{k}-(\hat{F}-F)^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F}\hat{\Gamma}^{-1}]^{-1}=I_{k}+(\hat{F}-F)^{\prime}V_{\varepsilon}^{-1}F\Gamma^{-1}+o_{p}(\frac{1}{\sqrt{n}})$. By plugging into (B.8), we get:
$\hat{F}-F+F[(\hat{F}-F)^{\prime}V_{\varepsilon}^{-1}F\Gamma^{-1}]=\frac{1}{\sqrt{n}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+o_{p}(\frac{1}{\sqrt{n}}).$ (B.9)
@@
By multiplying both sides with $M_{F,V_{\varepsilon}}$, we get $M_{F,V_{\varepsilon}}(\hat{F}-F)=\frac{1}{\sqrt{n}}M_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+o_{p}(\frac{1}{\sqrt{n}})$. Then, $\hat{F}-F=\frac{1}{\sqrt{n}}M_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+\frac{1}{\sqrt{n}}FA+o_{p}(\frac{1}{\sqrt{n}})$, where $A$ is a random $k\times k$ matrix to be determined next. By plugging into (B.9), we get $F(A+A^{\prime})=P_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+o_{p}(\frac{1}{\sqrt{n}})$. By multiplying both sides by $\frac{1}{2}\Gamma^{-1}F^{\prime}V_{\varepsilon}^{-1}$ and using $F^{\prime}V_{\varepsilon}^{-1}P_{F,V_{\varepsilon}}=F^{\prime}V_{\varepsilon}^{-1}$, we get the symmetric part of matrix $A$, i.e., $\frac{1}{2}(A+A^{\prime})=\frac{1}{2}\Gamma^{-1}F^{\prime}V_{\varepsilon}^{-1}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}$ (we include higher-order terms in the remainder $o_{p}(\frac{1}{\sqrt{n}})$). Thus, $\hat{F}-F=\frac{1}{\sqrt{n}}\Psi_{F}+o_{p}(\frac{1}{\sqrt{n}})$, where
$\Psi_{F}=M_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+\frac{1}{2}P_{F,V_{\varepsilon}}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F\Gamma^{-1}+F\tilde{A}$, (B.10)